OpenStudy (anonymous):

12^2-28x=-15 using the factoring method, solve for the roots of the quadratic equation. Be sure to place the equation in standard form before factoring. Check your answer.

OpenStudy (mathmale):

12^2-28x=-15 is incomplete. Something's missing. Please fix it.

OpenStudy (anonymous):

12x^2-28x+15=0

OpenStudy (anonymous):

it is complete

OpenStudy (anonymous):

12x^2-18x-10x+15=0 6x(2x-3)-5(2x-3)=0 (2x-3)(6x-5)=0

OpenStudy (anonymous):

x=3/2 or x=5/6

OpenStudy (anonymous):

it was both

OpenStudy (mathmale):

@candyhuebner : Was your "it was both" a statement or a question? What kind of response do you want to that statement?